OpenStudy (anonymous):

How to solve -(3-a)^2+4(5-a)?

OpenStudy (anonymous):

mostly the 1st part

OpenStudy (anonymous):

What are you trying to solve here? Is this the entire problem?

OpenStudy (cgreenwade2000):

If this is an equation, you need to have an equal sign somewhere.

OpenStudy (anonymous):

I don't get how to do the part when theres an exponent outside the equation of 2

OpenStudy (anonymous):

Just to simplyfy

OpenStudy (anonymous):

Oh, you need to use the formula (a+b)^2 = a^2 + 2ab + b^2. Replace the values you have and see if you can get the result.

OpenStudy (anonymous):

I don't understand what that is? Isn't the 2 just put into the 3 which is 9 and a^2

OpenStudy (anonymous):

why 2ab?

OpenStudy (anonymous):

Because you need to account for what is inside the parentheses. Here is an example: (2+3)^2 = 2^2 + 2*2*3 + 3^2 = 4 + 12 + 9 = 25 We know that 2+3=5, therefore 5^2 = 25; Now do this with your numbers. And take care with the minus sign.

OpenStudy (anonymous):

I get the part with 2^2 and 3^2 but i didnt know you had to add them inside and square them too

OpenStudy (anonymous):

So what would you simplifie the equation at the top

OpenStudy (anonymous):

Your correct result would be -(3-a)^2 = -(3^2 - 2*3*a + a^2) = -9 + 6a - a^2

OpenStudy (anonymous):

k ty